8. find the equation of the the line passing thru (4,-6) and (-3,5)
9. find the equation of the line with a slope of 0 and passing through the point (8,-6)

Answer by TimothyLamb(4379)

(4,-6) and (-3,5)
slope = m = dy/dx
m = (5+6)/(-3-4)
m = (-11/7)
y - 5 = (-11/7)(x + 3)
y = (-11/7)x + 3*(-11/7) + 5
y = (-11/7)x - 33/7 + 35/7
y = (-11/7)x + 2/7
---
(8,-6)
slope = 0 = horizontal line
y = -6
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php